### v1

# Instructions for *ACL Proceedings

Abstract
LLM-as-a-Judge refers to the automatic modeling of preferences for responses generated by Large Language Models (LLMs), which is of significant importance for both LLM evaluation and reward modeling. Although generative LLMs have made substantial progress in various tasks, their performance as LLM-Judge still falls short of expectations. In this work, we propose Think-J, which improves generative LLM-as-a-Judge by learning how to think. We first utilized a small amount of curated data to develop the model with initial judgment thinking capabilities. Subsequently, we optimize the judgment thinking traces based on reinforcement learning (RL). We propose two methods for judgment thinking optimization, based on offline and online RL, respectively. The offline method requires training a critic model to construct positive and negative examples for learning. The online method defines rule-based reward as feedback for optimization. Experimental results showed that our approach can significantly enhance the evaluation capability of generative LLM-Judge, surpassing both generative and classifier-based LLM-Judge without requiring extra human annotations [^1] Codes and data are openly available at https://github.com/HuihuiChyan/Think-J.
Think-J: Learning to Think for Generative LLM-as-a-Judge
[…]
footnotetext: * The first three authors contributed equally.
1 Introduction
As the capabilities of generative LLMs continue to advance, accurately evaluating the response quality has emerged as a crucial challenge. This is not only vital for more efficient model development and comparison but also essential in the context of Reinforcement Learning from Human Feedback (RLHF), which relies on precise preference modeling as guidance Wang et al. (2024a). However, traditional evaluation methods for generative models, such as BLEU Papineni et al. (2002), are based on predefined reference answers, which are often unavailable in open-ended scenarios.
Figure 1: Comparison of different judge models. Our proposed Think-J takes into account both accuracy and interpretability based on thinking optimization.
Some studies have proposed LLM-as-a-Judge Zheng et al. (2023), which leverages the generative capabilities of LLMs for evaluating response quality. These work either directly leverage proprietary LLMs or fine-tune a smaller judge based on preference data Gu et al. (2024). However, the accuracy of their judgments remains unsatisfactory as revealed by recent benchmarks Lambert et al. (2024). Other research has suggested fine-tuning a classifier based on preference data Liu et al. (2024a). While this method can achieve higher judgment accuracy, it lacks interpretability due to its scalar output, and the performance is highly dependent on the data quality Wang et al. (2024b).
Inspired by recent reasoning models such as o1 Jaech et al. (2024) and Deepseek-R1 Guo et al. (2025), in this work, we propose Thinking-enhanced Generative Judge (Think-J). Think-J aims to train a better generative judge by optimizing the model’s judgment thinking capabilities. Specifically, our method consists of two steps:
1) Judgment Thinking Initialization. We carefully curated 707 samples from preference data, considering various aspects such as accuracy, difficulty, and diversity. After that, thinking trace is annotated by proprietary models to initialize the thinking capabilities of the judge.
2) Judgment Thinking Optimization. Due to the lack of high-quality critique annotation in preference datasets, we opt to optimize judgment thinking ability based on reinforcement learning (RL). Specifically, we adopted two methods: a) Critic-guide Offline Learning, leveraging an additional critic model to generate corresponding thinking traces based on provided judgment results, thus constructing positive and negative examples for offline RL. b) Rule-based Online Learning, defining rule-based rewards based on the correctness of judgment results, thus optimizing the thinking trace by online RL Shao et al. (2024).
We conducted experiments on three open-source models, and results showed that our proposed Think-J significantly outperformed existing LLM-judges with only limited training data. We also verify the effectiveness of our method compared with generative and classifier-based preference modeling methods. Our contributions are as follows:
  1. 1.
    We propose to stimulate the judgment thinking ability of generative models with a small amount of carefully curated data.
  2. 2.
    We propose to optimize the judgment thinking ability of generative models with the correctness of judgment results as feedback.
  3. 3.
    Our proposed Think-J significantly outperforms previous generative and classifier-based LLM-judges for preference modeling.
2 Related Work
After the emergence of LLMs, numerous efforts have been made to design a more effective method for LLM evaluation Chang et al. (2023). One of the most scalable and effective method is LLM-as-a-Judge Li et al. (2023b); Zheng et al. (2023), namely utilizing proprietary LLMs, especially GPT4 Achiam et al. (2023), to evaluate the LLM’s response. For example, AlpacaEval Li et al. (2023b) used the win rate compared with baseline response determined by GPT-4 as the evaluation result. MT-Bench Zheng et al. (2023) automatically scored the model’s answers using GPT-4 as the results. The GPT-4-based evaluator is proven to presents comparable or even better consistency and accuracy compared with human.
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. Therefore, follow-up works suggest fine-tuning language models locally for evaluations, including JudgeLM Zhu et al. (2023b), Auto-J Li et al. (2023a), Prometheus Kim et al. (2023), Prometheus-2 Zhu et al. (2023a), OffsetBias Park et al. (2024), etc. These work typically construct preference data with judgment annotations and then finetune open-sourced LLMs to generate the judgment. Despite these fine-tuned judge models all achieve comparable accuracy with proprietary models, the evaluation is mostly conducted on the in-domain testsets, and these works are verified with a low scalability on more general benchmarks Huang et al. (2024).
Another group of work fine-tunes a classifier on preference data based on the Bradley-Terry model, which is more commonly used on reward modeling. This approach is simple yet effective, as demonstrated on RewardBench Lambert et al. (2024) where most top-performing models are trained in a classification style Liu et al. (2024a). However, this method does not fully leverage the generative capabilities of LLMs, and is unable to provide rationales for its judgments, which is crucial for scalable evaluation. While recent work has begun to leverage the generative abilities of LLMs to combine critiques for scalar reward prediction Ke et al. (2024); Ye et al. (2024), these critiques are often distilled from stronger proprietary models and hardly influence the final prediction Liu et al. (2025b). Effectively integrating the generative abilities of LLMs into evaluation remains an open challenge Chen et al. (2025a); Whitehouse et al. (2025); Chen et al. (2025b); Wang et al. (2025a).
3 Approach
3.1 Judgment Thinking Initialization
Recent studies have shown that LLMs inherently possess long chain-of-thought (CoT) reasoning capabilities, which can be activated with a small amount of data Muennighoff et al. (2025); Ye et al. (2025). In this work, we also curate high-quality preference data, LIMJ707, to initialize the thinking capability of the judge model. LIMJ707 are selected based on three principles:
  • Accuracy: The judgment annotation should be correct. We leverage the high-quality preference data Skywork-Preference-v0.2222huggingface.co/datasets/Skywork/Skywork-Reward-Preference-80K-v0.2, which has been carefully validated to ensure accuracy.
  • Difficulty: The sample should be sufficiently challenging. We apply the judge models to perform judgment for the sample three times, and select those samples where at least one judgment is failed, as these samples are likely more difficult and reflect the insufficiency of the judge.
  • Diversity: The instruction should encompass various types to enhance judgment thinking capabilities in different aspects. We represent the instructions with an embedding model and then merge samples that are too close to each other [^3] For more details please refer to Appendix B.1..
The data statistics during constructing LIMJ707 are shown in Table 1. Based on LIMJ707, we construct judgment thinking trace by Deepseek-R1. After that, the annotated samples are used to initialize the model with judgment thinking capability by Supervised Fine-tuning Ouyang et al. (2022).
     Skywork-Reward-Preference-v0.2 Data Num
       - initially selected 10K
       - filtered by difficulty 1496
       - filtered by diversity 870
       - filtered by accuracy 707
Table 1: Data statistics during constructing LIMJ707.
Despite its superior reasoning capability, the thinking trace generated by R1 are excessively long and could possibly cause difficulty for further optimization and increase computational overhead. Specifically, for our case, we hypothesize that general LLM judgment does not necessarily require a detailed, step-by-step reasoning process. On the contrary, identifying one key factor is sufficient to make correct judgment in most cases.
Therefore, we introduce Trace Clipping to enhance the effectiveness of thinking initialization [^4] An example is presented in Appendix B.1.. Specifically, we observed that the CoTs generated by R1 are composed of two parts: a lengthy reasoning process, and an explanation which summarizes the reasoning process. Therefore, we remove the first part and use the second part as the output trace. This not only facilitates subsequent optimization but also enhances the judgment readability.
3.2 Judgment Thinking Optimization
Figure 2: The illustration of our proposed framework. We begin by constructing high-quality judgment thinking traces using curated principles and proprietary thinking models. Based on this data, we initialize a judge model and a critic model, both equipped with judgment thinking capability. After that, we optimize the capability of the judge model through two methods: Critic-guided Offline Learning and Rule-based Online Learning, resulting in Think-J.
To further enhance the alignment between the judge and human preference, the initialized thinking trace should be further optimized on preference data. However, preference data typically only includes binary labels without thinking trace annotations. Therefore, we propose judgment thinking optimization based on reinforcement learning (RL). Specifically, we propose two optimization approaches based on offline and online RL methods, respectively, as detailed in the following sections.
3.2.1 Critic-guided Offline Learning
Offline RL methods represented by Direct Preference Optimization (DPO) Rafailov et al. (2023) has been widely applied to LLM pipelines due to their efficiency and simplicity Grattafiori et al. (2024). In this work, we also aim to optimize the thinking ability based on offline learning.
Due to the lack of golden thinking annotation in preference dataset, we propose to train an additional critic model [^5] Notice the critic model here differs from the critic model in traditional RL which is used for advantage estimation., to help to construct training samples. Both the critic and the judge models are trained on the same data (i.e., LIMJ707), with the following distinctions:
  • Judge Model: Given instruction-responses, it generates the thinking trace and judgment result.
  • Critic Model: Given instruction-responses and judgment result, it generates the thinking trace.
Based on the two models, we can perform thinking optimization with the following steps:
  1. 1.
    First, leverage the judge to evaluate the input to generate the thinking traces and results.
  2. 2.
    If the result is correct, use the critic to generate an incorrect trace as the negative sample. Conversely, if the result is incorrect, use the critic to generate a correct trace as the positive sample.
  3. 3.
    Based on the positive and negative samples, optimize the judgment thinking ability with offline learning objective.
  4. 4.
    These steps can be iterated to continuously enhance the judgment thinking capability.
We adopt a combination of SFT and DPO as our training objective in this step:
$\mathcal{L}_{\text{offline}}(\pi_{\theta};\mathcal{D})=$
$\>\>-\mathbb{E}_{x\sim\mathcal{D},(y_{w},y_{l})\sim\pi_{\theta}(y|x)}\left[\log\pi_{\theta}(y_{w}\mid x)+\right.$
$\>\>\left.\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{ref}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{ref}(y_{l}\mid x)}\right)\right]$ (1)
where $\pi_{\theta}$ and $\pi_{ref}$ represent the policy model and the reference model, and $y_{w}$ and $y_{l}$ denotes the positive and negative judgment traces, respectively.
With the help of critic model, samples with thinking annotation are constructed based on the correctness of judgment result. Therefore, the judge model will be enhanced to generate more accurate thinking, leading to more accurate judgment.
3.2.2 Rule-based Online Learning
The recent success of R1-style methods have demonstrated the effectiveness of online RL using discrete, rule-based rewards Shao et al. (2024). In this work, we also apply online rule-based RL approach to optimize the judgment thinking capability. More specifically, we mainly utilize the GRPO algorithm, whose optimization objective is:
$J_{\text{online}}(\pi_{\theta};\mathcal{D})=$
$\>\>\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(y|x)}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)}A_{i},\right.\right.$
$\>\>\left.\left.\text{clip}\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)},1-\epsilon,1+\epsilon\right)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\right]$ (2)
where $G$ is group size, and $A_{i}$ is advantage. The reward function is designed as follows [^6] We do not include a length penalty in rewards to encourage longer thinking, as we have observed that longer thinking does not necessarily lead to better accuracy in our case.:
$r_{\text{accuracy}}$ =$=\begin{cases}1,&\text{if }\text{judgement}=\text{label}\\ 0,&\text{if }\text{judgement}\neq\text{label}\end{cases}$ (3)
$r_{\text{format}}$ =$=\begin{cases}0,&\text{if }\text{format is right}\\ -0.5,&\text{if }\text{format is wrong}\end{cases}$ (4)
$r_{\text{strength}}$ $=||\text{s}_{\text{pred}}-\text{s}_{\text{golden}}||,s\in\{1,2,3\}$ (5)
$r_{\text{final}}$ $=\alpha\cdot r_{\text{accuracy}}+\beta\cdot r_{\text{format}}+\gamma\cdot r_{\text{strength}}$ (6)
Notice we incorporate a reward for assessing preference strength, which is defined as the degree to which the judge favors one response over another [^7] For example, a strength of 1 means the chosen response is only slightly better than the rejected, while a strength of 3 means the chosen is much better than the rejected.. The final judgment is also adjusted as:
<think> {thinking trace} </think>
Therefore, Response (a) is better, and the preference strength is [[2]].
Preference strength helps to perceive the relative quality of response pairs, without uniformly providing the same reward for different pairs despite the quality gap. We employ a comparatively simple scale of reward scores of reward, as during actual training, we observed that the model tends to manipulate scores towards extreme values  [^8] For more details please refer to Appendix A.1..
While annotating an absolute score for a given response is challenging, annotating preference strength is relatively easy, making strength annotation more readily accessible Wang et al. (2024d, 2025b). On the other hand, if absolute score annotation is absent in the preference data, we can firstly fine-tune a BT-classifier based on the data, then leverage the classifier to assess the relative strength between chosen and rejected responses, as explained in Wang et al. (2024b).
Model Sample Num RewardBench RMBench Think-J
Chat Hard Safety Reason Overall Overall Agreement
Claude-3-5-Sonnet-20240620999www.anthropic.com/news/claude-3-5-sonnet 96.4 74.0 81.6 84.7 84.2 68.9 70.7
Qwen-2.5-32B-Instruct 96.2 74.0 88.7 86.9 86.5 68.3 59.6
GPT-4o-2024-08-06101010platform.openai.com/docs/models/gpt-4o 96.1 76.1 88.1 86.6 86.7 68.8 69.8
Gemini-1.5-Pro-0514111111deepmind.google/technologies/gemini/pro 92.3 80.6 87.9 92.0 88.2 74.4 68.1
JudgeLM-33B Zhu et al. (2023b) 100K 90.1 51.0 85.7 39.7 66.6 49.6 45.3
Prometheus-7b-v2.0 Zhu et al. (2023a) 40K 83.9 49.2 72.8 72.0 69.5 52.4 63.1
Prometheus-8x7b-v2.0 Zhu et al. (2023a) 40K 93.0 47.1 80.5 77.4 74.5 57.4 68.5
Llama-3-OffsetBias-8B Park et al. (2024) 276K 92.5 80.3 86.8 76.4 84.0 66.0 68.7
CompassJudger-32B Cao et al. (2024) 2041K 97.4 65.6 85.1 87.1 83.8 69.4 80.7
STE-Llama3.1-70B Wang et al. (2024c) 20K 96.9 85.1 89.6 88.4 90.0 65.3 72.0
SynRM-Command-R-35B Ye et al. (2024) 5K 97.5 76.8 88.5 86.3 87.3
CLoud-Llama3-70B Ke et al. (2024) 350K 98.0 75.6 87.6 89.0 87.6
Think-J-Qwen-2.5-32B 9.8K 96.7 83.2 90.1 92.0 90.5 79.8 75.8
Table 2: Experiment results of different LLM judges on RewardBench, RMBench and Auto-J.
Model Method HH-RLHF HelpSteer2-Preference
RewardBench RewardBench
Chat Hard Safety Reason Overall Chat Hard Safety Reason Overall
Llama3-8B -Instruct Direct Prompt (w/o CoT) 90.4 44.7 76.5 63.4 68.8 90.4 44.7 76.5 63.4 68.8
Direct Prompt (w/ CoT) 84.6 40.9 50.8 58.8 58.8 84.6 40.9 50.8 58.8 58.8
SFT on LIMJ707 91.8 67.1 83.0 64.2 76.5 91.8 67.1 83.0 64.2 76.5
SFT (w/o CoT) 88.1 50.6 81.0 69.1 72.2 88.0 41.1 41.5 47.8 54.6
SFT (w/ CoT) 78.8 67.0 81.0 62.1 72.2 84.5 71.3 80.8 62.6 74.8
BT Classifier 81.8 72.6 79.3 85.2 79.7 88.6 73.9 81.2 91.8 83.9
CLoud 87.4 69.3 87.3 77.6 80.4 93.3 74.8 80.8 73.7 80.6
SynRM 87.2 74.3 81.1 80.2 80.7 91.9 68.4 80.4 87.9 82.2
Think-J 92.2 71.7 84.8 75.6 81.1 93.9 74.3 90.3 77.8 84.1
Qwen2.5-7B -Instruct Direct Prompt (w/o CoT) 94.4 56.9 81.0 77.3 77.4 94.4 56.9 81.0 77.3 77.4
Direct Prompt (w/ CoT) 93.0 58.1 81.2 77.8 77.5 93.0 58.1 81.2 77.8 77.5
SFT on LIMJ707 89.5 75.2 80.7 74.5 80.0 89.5 75.2 80.7 74.5 80.0
SFT (w/o CoT) 89.8 74.3 82.5 61.3 77.0 95.0 75.9 87.2 75.8 83.5
SFT (w/ CoT) 94.1 74.0 86.0 64.7 79.7 88.7 69.7 84.5 76.1 79.8
BT Classifier 82.4 69.7 87.0 76.9 79.0 95.0 67.8 85.0 61.4 77.3
CLoud 85.2 79.2 86.6 57.9 77.2 94.7 73.3 82.3 59.0 77.3
SynRM 90.5 65.1 77.8 70.6 76.0 96.7 61.4 82.8 61.4 75.6
Think-J 94.4 70.2 83.8 79.8 82.0 96.1 78.6 85.9 80.4 85.3
Table 3: Experiment results of different LLM judge training methods on RewardBench.
4 Experiments
4.1 Set-up
We mainly conducted experiments on two popular open-sourced models with their instruction version: Qwen-2.5 Qwen et al. (2025) and Llama-3 Grattafiori et al. (2024).
The optimization is conducted on two datasets: HelpSteer2-Preference121212huggingface.co/datasets/nvidia/HelpSteer2 and HH-RLHF131313huggingface.co/datasets/Anthropic/hh-rlhf. Notice while there are larger preference datasets available, we did not include them in our training, as our primary objective was to explore the most effective method for training LLM judges.
We compare Think-J with the following generative judgment approaches:
  • Direct Prompt Leverage the LLM to directly generate judgment.
  • SFT (w/o CoT) Train a generative model based on Supervised Fine-tuning to perform judgment without generating CoTs.
  • SFT (w/ CoT) Train a generative model on correct judgment thinking traces and results, which can also be deemed as Rejection Sampling.
We also compare Think-J with the following classifier-based approaches:
  • BT Classifier Feed the instruction and responses into the model and added a classification head, training it according to Bradley-Terry model.
  • CLoud Ke et al. (2024) First train the model to generate critiques, and then leverage the critiques as additional input to improve the BT Classifier.
  • SynRM Ye et al. (2024) Leverage critiques generated by a proprietary model as additional input to improve the BT Classifier. We use Deepseek-R1 to generate the critiques.
To further showcase Think-J’s evaluation capabilities, we also compared its 32B version against leading proprietary judge models, including closed-source options like Claude 3.5 Sonnet, GPT-4o, and Gemini 1.5 Pro, as well as open-source fine-tuned judges such as JudgeLM, Prometheus, and CompassJudger. These models are widely employed as LLM-as-a-Judge across diverse tasks.
We mainly perform evaluation on RewardBench Lambert et al. (2024), which is the most popular benchmark for evaluating preference modeling ability. We also evaluate our model on RMBench Liu et al. (2025a) and Auto-J-test Li et al. (2023a).
We report the best results achieved by either offline or online learning in the experiment results by default. We also mixed LIMJ707 into all training sets to ensure a fair comparison.
4.2 Main Experiment
As demonstrated in Table 2, Think-J achieved the best performance across all benchmarks, surpassing both close-sourced and fine-tuned judges. Notably, our method required only 9832 training samples, but still achieve marginal improvement on 32B-sized models. This underscores the effectiveness of Think-J, which leverages thinking optimization with the correctness of judgment as feedback to enhance preference modeling capability.
Furthermore, as Table 3 illustrates, starting from the same base model and training data, our proposed method consistently outperforms other approaches for fine-tuning LLM judges. This demonstrates the effectiveness of our judgment thinking optimization. In contrast, both naive and fine-tuned generative methods yield inferior results. Notably, the rejection sampling method, despite being trained on the same data constructed by the critic, also underperforms. This underscores the critical role of learning from negative samples when modeling human preference Liu et al. (2024b).
On the other hand, the classifier-based method achieves relatively higher results but can only produce numerical outputs that lack interpretability. Additionally, the reasoning-enhanced classifiers, including CLoud and SynRM, performs worse than expected. This suggests that combining generative CoT into classifier may introduce noise rather than useful information for classification.
5 Analysis
5.1 Less is More for Thinking Initialization
We compared the performance of different data source for judgment thinking initialization in Figure 3. Our findings indicate that a small amount of thinking trace annotated with Deepseek-R1 can significantly enhance the model’s judgment capabilities. In contrast, using thinking trace annotated with Deepseek-V3, or removing the trace from the data would results in a substantial decline in performance. This highlights the importance of high-quality trace for thinking initialization.
We also compare the impact of different data selection strategies in Table 5. The results show that data quality is crucial for effective model initialization. For instance, a model initialized with chatbot-arena141414huggingface.co/datasets/lmsys/chatbot_arena_conversations, which contains relatively noisy data, achieves minimal improvement. Conversely, selecting the longest traces proves detrimental, as the longest traces are often code or math-related, which can negatively impact the data diversity.
Llama-3-8B-Instruct Qwen-2.5-7B-Instruct
Method RewardBench RewardBench
Chat Hard Safety Reason Overall Chat Hard Safety Reason Overall
baseline 90.4 44.7 76.5 63.4 68.8 94.4 56.9 81.0 77.3 77.4
offline (Sampling) 92.9 75.0 86.9 70.1 81.2 92.9 72.4 85.3 76.2 81.7
offline (Critic-guided) 95.0 73.3 87.4 77.2 83.2 94.8 74.2 83.5 80.5 83.3
offline (w/o SFT) 95.1 63.8 88.3 77.6 81.2 95.7 73.8 86.0 74.6 82.5
online (PPO) 70.4 75.3 77.2 70.0 73.2 88.7 69.7 84.5 76.1 79.8
online (Reinforce++) 93.9 74.3 90.3 77.8 84.1 94.7 70.6 90.4 82.3 84.5
online (GRPO) 93.9 74.3 90.3 77.8 84.1 96.1 78.6 85.9 80.4 85.3
Table 4: Experiment results of different RL methods.
Figure 3: Experiment results of different data source for judgment thinking initialization.
Method RewardBench
Chat Safety Reason Overall
Llama-3-8B-Inst 64.8 76.5 63.4 68.8
Init with 1000 samples on chatbot-arena
random-sampled 66.0 74.3 64.4 68.3
Init with 1000 samples on skywork-preference-v3
random-sampled 75.1 82.0 66.4 75.4
longest 77.2 79.3 61.8 74.2
Llama3-failed 76.4 85.0 66.5 76.1
Table 5: Experiment results of different data selection strategies for judgment thinking initialization.
Init RL RewardBench
Chat Safety Reason Overall
Llama-3-8B-Inst 64.8 76.5 63.4 68.8
no init offline 80.3 73.2 79.4 79.1
online 80.8 86.8 71.8 80.8
w/o clip offline 82.6 88.1 70.2 81.4
online 79.6 85.1 67.5 78.7
w/ clip offline 82.8 87.4 77.2 83.2
online 82.9 90.3 77.8 84.1
Table 6: Experiment results of different initialization strategies for RL optimization on HelpSteer2.
We further investigated the impact of different initialization strategies on the subsequent optimization process, as shown in Table 6. The results indicate that even without thinking initialization, RL-based methods can still achieve substantial improvements, validating their effectiveness. Moreover, initializing the model with a few R1-annotated traces leads to a more structured and effective reasoning pattern, resulting in further enhancements in performance. In contrast, when initialization is performed without trace clipping, the optimization effect deteriorates, particularly for online RL. This finding supports our earlier hypothesis that for general LLM judgment, longer thinking do not necessarily correlate with better accuracy.
5.2 Best Practice for Thinking Optimization
Method RewardBench
Chat Safety Reason Overall
Llama-3-8B-Inst 64.8 76.5 63.4 68.8
Offline learning on Helpsteer2
iteration 1 81.8 88.3 77.8 83.2
iteration 2 82.5 87.2 74.8 82.5
iteration 3 82.8 87.4 77.2 83.2
Offline learning on HH-RLHF
iteration 1 81.2 83.2 67.0 77.3
iteration 2 81.1 84.6 70.5 78.4
iteration 3 78.3 83.1 72.6 78.9
Table 7: Experiment results of iterative offline learning.
Setting Qwen2.5-32B-Instruct Meta-Llama-3-8B-Instruct
RewardBench RewardBench
Chat Hard Safety Reason Overall Chat Hard Safety Reason Overall
$\beta$=1.0 96.7 83.2 90.1 92.0 90.5 93.9 74.3 90.3 77.8 84.1
$\beta$=0.5 96.9 85.8 91.2 87.4 90.3 95.0 73.3 87.4 77.2 83.2
$\beta$=0.0 92.9 71.3 86.3 67.3 79.4 96.2 65.5 88.0 71.5 80.3
$\gamma$=0.5 91.2 75.2 71.2 61.5 74.8 82.7 65.5 80.4 66.4 73.7
$\gamma$=0.2 96.7 83.2 90.1 92.0 90.5 96.5 66.8 88.7 73.2 81.3
$\gamma$=0.0 96.0 83.7 90.5 87.7 89.5 93.9 74.3 90.3 77.8 84.1
Table 8: Experiment results of different reward function settings on Helpsteer2-Preference.
With new RL algorithms continue to emerge in LLM training Zhang et al. (2025), in this section, we compare different RL methods and strategies for judgment thinking optimization.
As shown in Table 4, for offline RL, constructing traces based on sampling (with a large beam size) would lead to performance degradation. This verifies the critical role of critic model for constructing both positive and negative examples. On the other hand, for online RL methods, we find that PPO Schulman et al. (2017) significantly underperforms compared to GRPO and Reinforce++ Hu et al. (2025). This discrepancy suggests that incorporating a value model for thinking optimization would decrease training stability. We argue that natural language generation (NLG) tasks is distinct from the sequential decision-making tasks in traditional RL. Therefore, the introduction of an additional value model is not only unnecessary but may also hinder training efficiency.
Finally, perform offline learning in an iterative manner, as shown in Table 7, can achieve further improvement. However, this will result in a more complex and cumbersome training pipeline.
5.3 Reward Design for Online Learning
In this section, we aimi to analyze the contribution of different components of the reward function as defined in 6. We fixed $\alpha$ as 1.0 and varied the weights $\beta$, and $\gamma$, with the results detailed in Table 8. Our findings indicate that the formatting reward $r_{\text{format}}$ is crucial for both models, as its removal leads to significant performance degradation. Conversely, the influence of the strength reward $r_{\text{strength}}$ differs between the models. For Llama3-8B, removing $r_{\text{strength}}$ results in a performance improvement, while it offers a slight enhancement for the more capable Qwen2.5-32B. We hypothesize that Qwen2.5-32B’s stronger inherent abilities allow it to effectively learn the correlation between preference strength and judgment. In contrast, for the weaker model Llama-3-8B, this additional learning objective may introduce confusion.
5.4 Thinking Makes the Judge More Robust
In real applications, it is common that there exists noise in the training set, or there is a distributional difference between the training and test sets. In such cases, Think-J demonstrates superior robustness as a result of its judgment thinking ability.
To verify this, we adopted the approach from Wang et al. (2024b) and divided HH-RLHF into four groups with different data quality [^15] Data quality is indicated by the difference in scores assigned to response pairs by the reward model. For more details, please refer to Wang et al. (2024b).. We then trained judges on the data based on classifier-judge or Think-J. As shown in Figure 4, for the two groups with higher data quality, classifier-based judge achieve comparable or even better performance. However, for the groups with lower quality, the accuracy of classifier-based methods drops significantly, even falling below random guessing. In contrast, Think-J maintains relative stability, verifying its robustness to varied data quality.
Figure 4: Comparison of different methods on data groups with different quality. Group 1 is with lower quality and Group 4 is with higher quality.
6 Conclusion
In this paper, we propose Think-J to enhance generative LLM-as-Judge with judgment thinking optimization. Experiment results verify the effectiveness of Think-J compared with both classifier-based and generative LLM judges.
With the increasing popularity of RL-based test-time scaling methods, it is crucial to develop a reliable and stable feedback system that aligns well with real-world human preferences. In future, we will continue to explore generative judges for more accurate preference modeling and optimization.
Limitations
Our work still has several limitations: 1) Due to time and resource constraints, we only validated our method on 32B-sized models with a relatively small dataset. Fine-tuning larger models with more comprehensive preference data warrants future investigation. 2) While LLM-as-a-Judge and Reward Modeling are two inter-correlated tasks, the lack of absolute scoring data currently prevents us from extending our pairwise-trained LLM-Judge to a pointwise setting, hindering it for direct RL guidance. Acquiring such data is a key direction for future work. 3) The increasing prevalence of thinking-enhanced models like Qwen-3 necessitates investigating the continued effectiveness of our method on these architectures. Scaling our approach to reasoning-enhanced models would further strengthen our findings.
Ethical Considerations
Our research aims to model human preference, ultimately making LLM outputs less harmful and more helpful. While the judge model might occasionally produce relatively harmful outputs during experimentation, its core purpose is to provide unbiased judgment towards helpfulness and harmlessness, thereby reducing the occurrence of such outputs. Furthermore, our method leverages existing datasets rather than creating new ones. Based on these points, we believe our work is ethical.
Appendix A Additional Experiments
A.1 Margin-based Reward for GRPO
As we have explained in Section 3.2.2, we design the reward function as the combination of three parts: $r_{\text{accuracy}}$, $r_{\text{format}}$, $r_{\text{strength}}$. Referring to the work of Chen et al. (2025a), for the purpose of more fine-grained judgment prediction, we also tested the following reward function design:
$r_{\text{margin}}$ =$=\begin{cases}-||\text{s}_{\text{resp1}}-\text{s}_{\text{resp2}}||,&\text{if }\text{judgement}=\text{label}\\ ||\text{s}_{\text{resp1}}-\text{s}_{\text{resp2}}||,&\text{if }\text{judgement}\neq\text{label}\end{cases}$
where the judgment prediction should also be formatted as:
<think> {thinking trace} <think>
Therefore, the quality scores for Response (a) and Response (b) are [[30]] and [[50]], respectively.
Method Llama-3-8B-Instruct Qwen-2.5-7B-Instruct
RewardBench RewardBench
Chat Hard Safety Reason Average Chat Hard Safety Reason Average
online (GRPO w/ margin) 96.5 66.8 88.7 73.2 81.3 96.1 72.3 83.4 76.1 82.0
online (GRPO w/o margin) 93.9 74.3 90.3 77.8 84.1 96.1 78.6 85.9 80.4 85.3
Table 9: Experiment results of the impact of $r_{margin}$ on Helpsteer2-Preference.
This design enables fine-grained absolute response scoring, which is more useful for LLM preference optimization pipelines. However, as shown in Table 9, the introduction of margin-based reward results in performance degradation. By further inspection, we revealed a consistent issue of reward hacking, where the assigned scores were always driven to extreme values (e.g., 0 or 100) to maximize the reward. Despite experimenting with various formulations of $r_{\text{margin}}$, we were unable to effectively mitigate this problem. We believe that mitigating this reward hacking for $r_{\text{margin}}$ necessitates preference data annotated with absolute scores, which we leave for future investigation.
A.2 Statistical Indicators of Judgment Thinking Optimization
In this section, we present the detailed statistical indicators during Judgment Thinking Optimization. The statistical indicators of offline learning, GRPO training, PPO Schulman et al. (2017) training and Reinforce++ Hu et al. (2025) training are presented in Figure 5, 6, 8, 7, respectively.
As the results indicate, GRPO demonstrates consistent and substantial increases in predicted rewards, a strong signal of learning to generate desirable outputs, and the critic’s low and stable value function loss suggests reliable outcome evaluation. While DPO effectively aligns with preferences, and Reinforce++ and PPO show progress in reward acquisition and stability, GRPO’s robust reward improvement combined with a well-functioning critic positions it as the most robust algorithm based on these indicators.
Notably, GRPO’s response length exhibits a consistent gradual increase, similar to prior work of Shao et al. (2024), while other methods start to fluctuate after a certain point. This suggests GRPO effectively learns to generate more detailed reasoning, enhancing judgment accuracy.
Figure 5: The variation of statistical metrics during offline learning trained on Helpsteer2-Preference.
Figure 6: The variation of statistical metrics during GRPO training on Helpsteer2-Preference.
Figure 7: The variation of statistical metrics during Reinforce++ training on Helpsteer2-Preference.
Figure 8: The variation of statistical metrics during PPO training on Helpsteer2-Preference.
Appendix B Implementation Details
B.1 Judgment Thinking Initialization
As we have explained in Section 3.1, we apply deduplication to ensure the diversity of LIMJ707. For this purpose, we leverage the algorithm defined in Algorithm 1 to create instruction data pool, where the embeddings of instructions are derived based on the model of instructor-large [^16] https://huggingface.co/hkunlp/instructor-large, and repetitive instructions are removed based on the similarity of instruction semantics.
We leverage the prompt templates in Figure 11 and 13 for constructing thinking trace based on proprietary thinking models. The temperature is set as 0.0 to enable more accurate judgment. We also present an example of thinking trace calibration in 10, which can ensure training efficiency and stability.
Algorithm 1 Diversity-verified Sampling
0:  Instruction Dataset $D$ with $m$ samples, number of samples to select $n$.
0:  Selected Dataset $D^{\prime}$ with $n$ samples.
1:  Derive the embeddings for each sample in $D$.
2:  Random sample one data point $x$ from $D$.
3:  Delete $x$ from $D$, add $x$ to $D^{\prime}$.
4:  for $i=1,2,...,n-1$ do
5:     Calculate the cosine similarity score between $x$ and each sample from $D$.
6:     Select the least similar sample $x^{\prime}$ from $D^{\prime}$.
7:     Let $x=$ $x^{\prime}$.
8:     Delete $x$ from $D$, add $x$ to $D^{\prime}$.
9:  end for
Figure 9: The sampling algorithm for LIMJ707 used for diversity verification.
Figure 10: The illustration of Trace Clipping. The first part of thinking trace is removed, while the second part is used as the output thinking trace.
B.2 Judgment Thinking Optimization
For judgment thinking optimization, we leverage the prompts in Figure 11 and 12 for offline learning, and the prompts in Figure 13 and 14 for online learning. The offline learning is implemented based on Llama-Factory Zheng et al. (2024), and online learning is implemented based on verl Sheng et al. (2024). We present the hyper-parameter settings in Table 10 and 11.
hyper-parameter SFT offline BT classifier
learning_rate 5.00E-06 1.00E-06 2.00E-06
per_device_train_batch_size 64 64 64
cutoff_len 4096 4096 4096
lr_scheduler cosine cosine cosine
train_epoch 3 2 2
kl_beta 0.1
deepspeed_stage zero3 zero2
precision bf16 bf16 bf16
flash_attn fa2 fa2
Table 10: Hyper-parameter settings for SFT, offline learning and BT classifier.
hyper-parameter GRPO PPO Reinforce++
data.train_batch_size 1024 1024 1024
data.max_prompt_length 1024 1024 1024
data.max_response_length 3072 3072 3072
actor_rollout_ref.actor.optim.lr 1.00E-06 1.00E-06 3.00E-06
actor_rollout_ref.actor.ppo_mini_batch_size 256 256 1024
actor_rollout_ref.actor.ppo_micro_batch_size_per_gpu 16 4 8
actor_rollout_ref.actor.use_kl_loss true true
actor_rollout_ref.actor.kl_loss_coef 0.001 0.001
actor_rollout_ref.actor.kl_loss_type low_var_kl mse
actor_rollout_ref.rollout.log_prob_micro_batch_size_per_gpu 16 4 8
actor_rollout_ref.rollout.n 8 4
actor_rollout_ref.ref.log_prob_micro_batch_size_per_gpu 16 4 8
critic.optim.lr 1.00E-05
critic.ppo_micro_batch_size_per_gpu 4
algorithm.kl_ctrl.kl_coef 0.001 0.001
trainer.n_gpus_per_node 8 8 8
trainer.nnodes 2 2 2
trainer.total_epochs 10 5 5
Table 11: Hyper-parameter settings for online learning algorithms.
Figure 11: The prompt template used for judgment prediction without $r_{strength}$.
Figure 12: The prompt template used for critique prediction without $r_{strength}$.
Figure 13: The prompt template used for judgment prediction with $r_{strength}$.
Figure 14: The prompt template used for critique prediction with $r_{strength}$.