### v2

# Think-J: Learning to Think for Generative LLM-as-a-Judge

Abstract
LLM-as-a-Judge refers to the automatic modeling of preferences for responses generated by Large Language Models (LLMs), which is of significant importance for both LLM evaluation and reward modeling. Although generative LLMs have made substantial progress in various tasks, their performance as LLM-Judge still falls short of expectations. In this work, we propose Think-J, which improves generative LLM-as-a-Judge by learning how to think. We first utilized a small amount of curated data to develop the model with initial judgment thinking capabilities. Subsequently, we optimize the judgment thinking traces based on reinforcement learning (RL). We propose two methods for judgment thinking optimization, based on offline and online RL, respectively. The offline method requires training a critic model to construct positive and negative examples for learning. The online method defines rule-based reward as feedback for optimization. Experimental results showed that our approach can significantly enhance the evaluation capability of generative LLM-Judge, surpassing both generative and classifier-based LLM-Judge without requiring extra human annotations.
footnotetext: * Equal contribution.footnotetext: $\dagger$ Corresponding Author.
Codehttps://github.com/huihuichyan/think-j
1 Introduction
As the capabilities of generative LLMs continue to advance, accurately evaluating the response quality has emerged as a crucial challenge(Huang et al. 2025). This is not only vital for more efficient model development and comparison but also essential in the context of Reinforcement Learning from Human Feedback (RLHF), which relies on precise preference modeling as guidance (Wang et al. 2024a; He et al. 2025; Liu et al. 2025c). However, traditional evaluation methods for generative models, such as BLEU (Papineni et al. 2002), are based on predefined reference answers, which are often unavailable in open-ended scenarios.
Figure 1: Comparison of different judge models. Our proposed Think-J takes into account both accuracy and interpretability based on thinking optimization.
Some studies have proposed LLM-as-a-Judge (Zheng et al. 2023), which leverages the generative capabilities of LLMs for evaluating response quality. These work either directly leverage proprietary LLMs or fine-tune a smaller judge based on preference data (Gu et al. 2024). However, the accuracy of their judgments remains unsatisfactory as revealed by recent benchmarks (Lambert et al. 2024). Other research has suggested fine-tuning a classifier based on preference data (Liu et al. 2024a). While this method can achieve higher judgment accuracy, it lacks interpretability due to its scalar output, and the performance is highly dependent on the data quality (Wang et al. 2024b).
Inspired by recent reasoning models such as o1 (Jaech et al. 2024) and Deepseek-R1 (Guo et al. 2025), in this work, we propose Thinking-enhanced Generative Judge (Think-J), as shown in Figure 1. Think-J aims to train a better generative judge by optimizing the model’s judgment thinking capabilities. Specifically, Think-J consists of two steps:
1) Judgment Thinking Initialization. We carefully curated 707 samples from preference data, considering various aspects such as accuracy, difficulty, and diversity. After that, thinking trace is annotated by proprietary models to initialize the thinking capabilities of the judge.
2) Judgment Thinking Optimization. Due to the lack of high-quality critique annotation in preference datasets, we opt to optimize judgment thinking ability based on reinforcement learning (RL). Specifically, we adopted two methods: a) Critic-guided Offline Learning, leveraging an additional critic model to generate corresponding thinking traces based on provided judgment results, thus constructing positive and negative examples for offline RL. b) Rule-based Online Learning, defining rule-based rewards based on the correctness of judgment results, thus optimizing the thinking trace by online RL (Shao et al. 2024).
We conducted experiments on three open-source models, and results showed that our proposed Think-J significantly outperformed existing LLM-judges with only limited training data. We also verify the effectiveness of our method compared with generative and classifier-based preference modeling methods. Our contributions are as follows:
  1. 1.
    We propose to stimulate the judgment thinking ability of generative models with carefully curated data.
  2. 2.
    We propose to optimize the judgment thinking ability of generative models with reinforcement learning.
  3. 3.
    Our proposed Think-J significantly outperforms previous generative and classifier-based LLM-as-judges.
2 Background
After the emergence of LLMs, numerous efforts have been made to design a more effective method for LLM evaluation (Chang et al. 2023). One of the most scalable and effective methods is LLM-as-a-Judge (Li et al. 2023b; Zheng et al. 2023), namely utilizing proprietary LLMs, especially GPT4 (Achiam et al. 2023), to evaluate the LLM’s response. For example, AlpacaEval (Li et al. 2023b) used the win rate compared with baseline response determined by GPT-4 as the evaluation result. MT-Bench (Zheng et al. 2023) automatically scored the model’s answers using GPT-4 as the results. The GPT-4-based evaluator is proven to presents comparable or even better consistency compared with human.
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. Therefore, follow-up works suggest fine-tuning language models locally for evaluations, including JudgeLM (Zhu et al. 2023b), Auto-J (Li et al. 2023a), Prometheus (Kim et al. 2023), Prometheus-2 (Zhu et al. 2023a), OffsetBias (Park et al. 2024), etc. These work typically construct preference data with judgment annotations and then finetune open-sourced LLMs to generate the judgment. Despite these fine-tuned judge models all achieve comparable accuracy with proprietary models, the evaluation is mostly conducted on the in-domain testsets, and these works are verified with a low scalability on more general benchmarks (Huang et al. 2024).
Another group of work fine-tunes a classifier on preference data based on the Bradley-Terry model, which is more commonly used on reward modeling. This approach is simple yet effective, as demonstrated on RewardBench (Lambert et al. 2024) where most top-performing models are trained in a classification style (Liu et al. 2024a). However, this method does not fully leverage the generative capabilities of LLMs, and is unable to provide rationales for its judgments, which is crucial for scalable evaluation. While recent work has begun to leverage the generative abilities of LLMs to combine critiques for scalar reward prediction (Ke et al. 2024; Ye et al. 2024), these critiques are often distilled from stronger proprietary models and hardly influence the final prediction (Liu et al. 2025d). Effectively integrating the generative abilities of LLMs into evaluation remains an open challenge (Chen et al. 2025a; Whitehouse et al. 2025; Chen et al. 2025b; Wang et al. 2025a).
Figure 2: The illustration of our proposed framework. We begin by constructing high-quality judgment thinking traces using curated principles and proprietary thinking models. Based on this data, we initialize a judge model and a critic model, both equipped with judgment thinking capability. After that, we optimize the capability of the judge model through two methods: Critic-guided Offline Learning and Rule-based Online Learning, resulting in Think-J.
3 Methodology
3.1 Judgment Thinking Initialization
Recent studies have shown that LLMs inherently possess long chain-of-thought (CoT) reasoning capabilities, which can be activated with a small amount of data (Muennighoff et al. 2025; Ye et al. 2025; Liu et al. 2025a). In this work, we also curate high-quality preference data, LIMJ707, to initialize the thinking capability of the judge model. Specifically, LIMJ707 is selected based on three principles:
  • Accuracy: The judgment (preference) annotation should be correct. We leverage the high-quality preference data Skywork-Preference-v0.2111huggingface.co/datasets/Skywork/Skywork-Reward-Preference-80K-v0.2, which has been carefully validated to ensure accurate annotation.
  • Difficulty: The sample should be sufficiently challenging. We apply the judge models to perform judgment for the sample three times, and select those samples where at least one judgment is failed, as these samples are likely more difficult and reflect the insufficiency of the judge.
  • Diversity: The instruction should encompass various types to enhance judgment thinking capabilities in different aspects. We represent the instructions with an embedding model and then merge duplicate samples [^2] For more details please refer to Appendix B.1..
The data statistics during processing are shown in Table 1. Based on LIMJ707, we construct judgment thinking trace by Deepseek-R1. After that, the annotated samples are used to initialize the model with judgment thinking capability by Supervised Fine-tuning (Ouyang et al. 2022) [^3] Due to the overly long thinking trace generated by Deepseek-R1, we performed trace-clipping to reduce training overhead and improve efficiency. Please refer to Appendix B.1 for more details..
   Skywork-Reward-Preference-v0.2 Data Num
     - initially selected 10K
     - filtered by difficulty 1496
     - filtered by diversity 870
     - filtered by accuracy 707
Table 1: Data statistics during constructing LIMJ707.
3.2 Judgment Thinking Optimization
To further enhance the alignment between the judge and human preference, the initialized thinking trace should be further optimized on preference data. However, preference data typically only includes binary labels without thinking trace annotations. Therefore, we propose judgment thinking optimization based on reinforcement learning (RL). Specifically, we propose two methods based on offline and online learning respectively, as shown in Figure 2.
Critic-guided Offline Learning
Offline RL methods represented by Direct Preference Optimization (DPO) (Rafailov et al. 2023) has been widely applied to LLM pipelines due to their efficiency and simplicity (Grattafiori et al. 2024). In this work, we also aim to optimize the judgment thinking ability based on offline learning.
Due to the lack of golden thinking annotation in preference dataset [^4] The rationale for fine-tuning an additional critic model over sampling to generate preference data is presented in Appendix A.3., we propose to train an additional critic model [^5] Notice the critic model here differs from the critic model in traditional RL which is used for advantage estimation., to help to construct training samples. Both the critic and the judge models are trained on the same data (i.e., LIMJ707), with the following distinctions:
  • Judge Model: Given instruction-responses, it generates the thinking trace and judgment result.
  • Critic Model: Given instruction-responses and judgment result, it generates the thinking trace.
Based on the two models, we can perform thinking optimization with the following steps:
  1. 1.
    First, leverage the judge to evaluate the input to generate the thinking traces and results.
  2. 2.
    If the result is correct, use the critic to generate an incorrect trace as the negative sample. Conversely, if the result is incorrect, use the critic to generate a correct trace as the positive sample.
  3. 3.
    Based on the positive and negative samples, optimize the judgment thinking ability with offline learning objective.
  4. 4.
    These steps can be iterated to continuously enhance the judgment thinking capability.
We adopt a combination of SFT and DPO as our training objective in this step:
$\mathcal{L}_{\text{offline}}(\pi_{\theta};\mathcal{D})=$
$\>\>-\mathbb{E}_{x\sim\mathcal{D},(y_{w},y_{l})\sim\pi_{\theta}(y|x)}\left[\log\pi_{\theta}(y_{w}\mid x)+\right$.
$\>\>\left.\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{ref}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{ref}(y_{l}\mid x)}\right)\right]$ (1)
where $\pi_{\theta}$ and $\pi_{ref}$ represent the policy model and the reference model, and $y_{w}$ and $y_{l}$ denotes the positive and negative judgment traces, respectively.
With the help of critic model, samples with thinking annotation are constructed based on the correctness of judgment result. Therefore, the judge model will be enhanced to generate more accurate thinking for better judgment.
Rule-based Online Learning
The recent success of R1-style methods have demonstrated the effectiveness of online RL using discrete, rule-based rewards (Shao et al. 2024). In this work, we also apply online rule-based RL approach to optimize the judgment thinking capability. More specifically, we mainly utilize the GRPO algorithm, with the optimization objective as follows:
$J_{\text{online}}(\pi_{\theta};\mathcal{D})=$
[…]
$\>\>\left.\left.\text{clip}\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)},1-\epsilon,1+\epsilon\right)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\right]$ (2)
where $G$ is group size, and $A_{i}$ is advantage. The reward function is designed as follows [^6] We do not include a length penalty in rewards to encourage longer thinking, as we have observed that longer thinking does not necessarily lead to better accuracy in our case.:
$r_{\text{accuracy}}$ $=\begin{cases}1,&\text{if }\text{judgement}=\text{label}\\ 0,&\text{if }\text{judgement}\neq\text{label}\end{cases}$ (3)
$r_{\text{format}}$ $=\begin{cases}0,&\text{if }\text{format is right}\\ -0.5,&\text{if }\text{format is wrong}\end{cases}$ (4)
[…]
$r_{\text{final}}$ $=\alpha\cdot r_{\text{accuracy}}+\beta\cdot r_{\text{format}}+\gamma\cdot r_{\text{strength}}$ (6)
where detailed weights for different rewards are presented in Section 5.2. Notice we incorporate a reward for assessing preference strength, which is defined as the degree to which the judge favors one response over another [^7] For example, a strength of 1 means the chosen response is only slightly better than the rejected, while a strength of 3 means the chosen is much better than the rejected. For more details about the criteria for the strength annotation, please refer to Appendix B.4.. The prompt template of judgment is also adjusted as:
<think> {thinking trace} </think>
Therefore, Response (a) is better, and the preference strength is [[2]].
Preference strength helps to perceive the relative quality of response pairs, without uniformly providing the same reward for different pairs despite the quality gap. We employ a comparatively simple scale of reward scores of reward, as during actual training, we observed that the model tends to manipulate scores towards extreme values [^8] Please refer to Appendix A.1 for more details ..
While absolute score annotation for a given response is challenging, annotating relative preference strength is more readily accessible (Wang et al. 2024d, 2025b). Moreover, if absolute score annotation is absent in the preference data, we can firstly fine-tune a BT-classifier based on the data, then leverage the classifier to assess the relative strength between chosen and rejected responses (Wang et al. 2024b).
Model Sample Num RewardBench RMBench Auto-J
Chat Hard Safety Reason Overall Overall Agreement
Claude-3-5-Sonnet-20240620 96.4 74.0 81.6 84.7 84.2 68.9 70.7
Qwen-2.5-32B-Instruct 96.2 74.0 88.7 86.9 86.5 68.3 59.6
GPT-4o-2024-08-06 96.1 76.1 88.1 86.6 86.7 68.8 69.8
Gemini-1.5-Pro-0514 92.3 80.6 87.9 92.0 88.2 74.4 68.1
JudgeLM-33B (Zhu et al. 2023b) 100K 90.1 51.0 85.7 39.7 66.6 49.6 45.3
Prometheus-7b-v2.0 (Zhu et al. 2023a) 40K 83.9 49.2 72.8 72.0 69.5 52.4 63.1
Prometheus-8x7b-v2.0 (Zhu et al. 2023a) 40K 93.0 47.1 80.5 77.4 74.5 57.4 68.5
Llama-3-OffsetBias-8B (Park et al. 2024) 276K 92.5 80.3 86.8 76.4 84.0 66.0 68.7
CompassJudger-32B (Cao et al. 2024) 2041K 97.4 65.6 85.1 87.1 83.8 69.4 80.7
STE-Llama3.1-70B (Wang et al. 2024c) 20K 96.9 85.1 89.6 88.4 90.0 65.3 72.0
SynRM-Command-R-35B (Ye et al. 2024) 5K 97.5 76.8 88.5 86.3 87.3
CLoud-Llama3-70B (Ke et al. 2024) 350K 98.0 75.6 87.6 89.0 87.6
Think-J-Qwen-2.5-32B (Helpsteer2-Pref) 9.8K 96.7 83.2 90.1 92.0 90.5 79.8 75.8
Table 2: Experiment results of top LLM-Judges on RewardBench, RMBench, and Auto-J-test. In this table, we report the best performance achieved by various LLM-Judges, trained on different base models and datasets.
Model Method $H$-RLHF Helpsteer2-Pref
RewardBench RewardBench
Chat Hard Safety Reason Overall Chat Hard Safety Reason Overall
Llama3-8B -Instruct Direct Prompt (w/o CoT) 90.4 44.7 76.5 63.4 68.8 90.4 44.7 76.5 63.4 68.8
Direct Prompt (w/ CoT) 84.6 40.9 50.8 58.8 58.8 84.6 40.9 50.8 58.8 58.8
SFT on LIMJ707 91.8 67.1 83.0 64.2 76.5 91.8 67.1 83.0 64.2 76.5
SFT (w/o CoT) 88.1 50.6 81.0 69.1 72.2 88.0 41.1 41.5 47.8 54.6
SFT (w/ CoT) 78.8 67.0 81.0 62.1 72.2 84.5 71.3 80.8 62.6 74.8
BT Classifier 81.8 72.6 79.3 85.2 79.7 88.6 73.9 81.2 91.8 83.9
CLoud 87.4 69.3 87.3 77.6 80.4 93.3 74.8 80.8 73.7 80.6
SynRM 87.2 74.3 81.1 80.2 80.7 91.9 68.4 80.4 87.9 82.2
Think-J 92.2 71.7 84.8 75.6 81.1 93.9 74.3 90.3 77.8 84.1
Qwen2.5-7B -Instruct Direct Prompt (w/o CoT) 94.4 56.9 81.0 77.3 77.4 94.4 56.9 81.0 77.3 77.4
Direct Prompt (w/ CoT) 93.0 58.1 81.2 77.8 77.5 93.0 58.1 81.2 77.8 77.5
SFT on LIMJ707 89.5 75.2 80.7 74.5 80.0 89.5 75.2 80.7 74.5 80.0
SFT (w/o CoT) 89.8 74.3 82.5 61.3 77.0 95.0 75.9 87.2 75.8 83.5
SFT (w/ CoT) 94.1 74.0 86.0 64.7 79.7 88.7 69.7 84.5 76.1 79.8
BT Classifier 82.4 69.7 87.0 76.9 79.0 95.0 67.8 85.0 61.4 77.3
CLoud 85.2 79.2 86.6 57.9 77.2 94.7 73.3 82.3 59.0 77.3
SynRM 90.5 65.1 77.8 70.6 76.0 96.7 61.4 82.8 61.4 75.6
Think-J 94.4 70.2 83.8 79.8 82.0 96.1 78.6 85.9 80.4 85.3
Table 3: Experiment results of different LLM judge training methods on RewardBench. In this table, we report the performance of various methods trained on the same base models (Llama3-8B-Instruct and Qwen2.5-7B-Instruct) and datasets ($H$-RLHF and Helpsteer2-Pref) to enable a more direct comparison.
4 Experiments
4.1 Set-up
We mainly conducted experiments on two popular open-sourced models with their instruction version: Qwen-2.5 (Qwen et al. 2025) and Llama-3 (Grattafiori et al. 2024).
We primarily conducted our optimization on two datasets: HelpSteer2-Pref999huggingface.co/datasets/nvidia/HelpSteer2 and HH-RLHF101010huggingface.co/datasets/Anthropic/hh-rlhf. For a fair comparison, LIMJ707 was mixed into all training sets. While larger preference datasets are available, we excluded them from our training because our primary objective was to explore the most effective method for training LLM judges.
We mainly compare Think-J with the following generative LLM-judge approaches:
  • Direct Prompt Leverage the LLM to directly generate judgment without fine-tuning.
  • SFT (w/o CoT) Train a generative model by supervised fine-tuning to perform judgment without thinking traces.
  • SFT (w/ CoT) Train a generative model on correct judgment thinking traces and results.
We also compare Think-J with the following classifier-based LLM-judge approaches:
  • BT Classifier Feed the instruction and responses into the model and added a classification head, training it according to Bradley-Terry model (Sun et al. 2025).
  • CLoud (Ke et al. 2024) First train the model to generate critiques, and then leverage the critiques as additional input to improve the BT Classifier.
  • SynRM (Ye et al. 2024) Leverage critiques generated by a proprietary model as additional input to improve the BT Classifier. We use Deepseek-R1 to generate the critiques.
To further showcase Think-J’s evaluation capabilities, we also compared its 32B version against leading proprietary judge models, including closed-source options like Claude 3.5 Sonnet, GPT-4o, and Gemini 1.5 Pro, as well as open-source fine-tuned judges such as JudgeLM, Prometheus, and CompassJudger. These models are widely employed as LLM-as-a-Judge across diverse tasks.
We mainly perform evaluation on RewardBench (Lambert et al. 2024). We also evaluate our model on RMBench (Liu et al. 2025b) and Auto-J-test (Li et al. 2023a).
We report the best results achieved by either offline or online learning for the main experiments by default.
4.2 Main Experiment
Llama-3-8B-Instruct Qwen-2.5-7B-Instruct
Method RewardBench RewardBench
Chat Hard Safety Reason Overall Chat Hard Safety Reason Overall
baseline 90.4 44.7 76.5 63.4 68.8 94.4 56.9 81.0 77.3 77.4
offline (Critic-guided) 95.0 73.3 87.4 77.2 83.2 94.8 74.2 83.5 80.5 83.3
offline (w/o SFT) 95.1 63.8 88.3 77.6 81.2 95.7 73.8 86.0 74.6 82.5
online (PPO) 70.4 75.3 77.2 70.0 73.2 88.7 69.7 84.5 76.1 79.8
online (Reinforce++) 93.7 74.3 89.7 77.9 84.0 94.7 70.6 90.4 82.3 84.5
online (GRPO) 93.9 74.3 90.3 77.8 84.1 96.1 78.6 85.9 80.4 85.3
Table 4: Experiment results of different RL strategies.
As demonstrated in Table 2, Think-J-32B achieved the best performance across all benchmarks, surpassing both close-sourced and fine-tuned judges. Notably, our method required only 9832 training samples, but still achieve marginal improvement on 32B-sized models. This underscores the effectiveness of Think-J, which leverages thinking optimization with the correctness of judgment as feedback to enhance preference modeling capability.
Furthermore, as Table 3 illustrates, starting from the same base model and training data, our proposed method consistently outperforms other approaches for fine-tuning LLM judges. This demonstrates the effectiveness of our judgment thinking optimization. In contrast, both naive and fine-tuned generative methods yield inferior results. Notably, the rejection sampling method, despite being trained on the same data constructed by the critic, also underperforms. This underscores the critical role of learning from negative samples when modeling human preference (Liu et al. 2024b).
On the other hand, the classifier-based method achieves relatively higher results but can only produce numerical outputs that lack interpretability. Additionally, the reasoning-enhanced classifiers, including CLoud and SynRM, performs worse than expected. This suggests that combining generative CoT into classifier may introduce noise rather than useful information for classification.
5 Analysis
5.1 Less is More for Thinking Initialization
We compared the performance of different data source for judgment thinking initialization in Figure 3. Our findings indicate that a small amount of thinking trace annotated with Deepseek-R1 can significantly enhance the model’s judgment capabilities. In contrast, using thinking trace annotated with Deepseek-V3, or removing the trace from the data would results in a substantial decline in performance. This highlights the importance of high-quality thinking trace for judgment thinking initialization.
We also compare the impact of different data selection strategies in Table 5. The results show that data quality is crucial for effective model initialization. For instance, a model initialized with chatbot-arena(Chiang et al. 2024), which contains relatively noisy data, achieves minimal improvement. Conversely, selecting the longest traces proves detrimental, as the longest traces are often code or math-related, which can negatively impact the data diversity.
Figure 3: Experiment results of different data source for judgment thinking initialization.
Method RewardBench
Chat Safety Reason Overall
Llama-3-8B-Inst 64.8 76.5 63.4 68.8
Init with 1000 samples on chatbot-arena
random-sampled 66.0 74.3 64.4 68.3
Init with 1000 samples on skywork-preference-v3
random-sampled 75.1 82.0 66.4 75.4
longest 77.2 79.3 61.8 74.2
Llama3-failed 76.4 85.0 66.5 76.1
Table 5: Experiment results of different data selection strategies for judgment thinking initialization.
Init RL RewardBench
Chat Safety Reason Overall
Llama-3-8B-Inst 64.8 76.5 63.4 68.8
no init offline 80.3 73.2 79.4 79.1
online 80.8 86.8 71.8 80.8
w/ init offline 82.8 87.4 77.2 83.2
online 82.9 90.3 77.8 84.1
Table 6: Experiment results of different initialization strategies for RL optimization on HelpSteer2-Pref.
Setting Qwen2.5-32B-Instruct Qwen2.5-7B-Instruct
RewardBench RewardBench
Chat Hard Safety Reason Overall Chat Hard Safety Reason Overall
$\beta$=1.0 96.7 83.2 90.1 92.0 90.5 96.1 78.6 85.9 80.4 85.3
$\beta$=0.5 96.9 85.8 91.2 87.4 90.3 95.0 77.3 87.6 77.2 84.3
$\beta$=0.0 92.9 71.3 86.3 67.3 79.4 95.2 71.5 88.0 71.7 81.6
$\gamma$=0.5 91.2 75.2 71.2 61.5 74.8 93.6 63.4 82.0 74.1 78.3
$\gamma$=0.2 96.7 83.2 90.1 92.0 90.5 94.7 76.3 88.7 78.2 84.5
$\gamma$=0.0 96.0 83.7 90.5 87.7 89.5 96.1 78.6 85.9 80.4 85.3
Table 7: Experiment results of different reward function settings on Helpsteer2-Pref.
Method RewardBench
Chat Safety Reason Overall
Llama-3-8B-Inst 64.8 76.5 63.4 68.8
Offline learning on Helpsteer2
iteration 1 81.8 88.3 77.8 83.2
iteration 2 82.5 87.2 74.8 82.5
iteration 3 82.8 87.4 77.2 83.2
Offline learning on HH-RLHF
iteration 1 81.2 83.2 67.0 77.3
iteration 2 81.1 84.6 70.5 78.4
iteration 3 78.3 83.1 72.6 78.9
Table 8: Experiment results of iterative offline learning.
We further investigated the impact of different initialization strategies on the subsequent optimization process, as shown in Table 6. The results indicate that even without thinking initialization, RL-based methods can still achieve substantial improvements, validating their effectiveness. Moreover, initializing the model with a few R1-annotated traces leads to a more structured and effective reasoning pattern, resulting in further enhancements in performance.
5.2 Best Practice for Thinking Optimization
With new RL algorithms continue to emerge in LLM training (Zhang et al. 2025), in this section, we compare different RL strategies for judgment thinking optimization.
As shown in Table 4, for offline learning, removing SFT term will lead to performance degradation, as SFT is conducted on token-level and can provide regularization for better preference optimization. On the other hand, for online learning, we find that PPO (Schulman et al. 2017) significantly underperforms compared to GRPO and Reinforce++ (Hu et al. 2025). This discrepancy suggests that incorporating a value model for thinking optimization would decrease training stability. We argue that natural language generation (NLG) tasks is distinct from the sequential decision-making tasks in traditional RL. Therefore, the introduction of an additional value model is not only unnecessary but may also hinder training efficiency.
Finally, as shown in Table 8, offline learning in an iterative manner can achieve further improvement. However, this will result in a more complex and cumbersome training pipeline. To draw a conclusion, the best practice for thinking optimization is GRPO with carefully designed rewards.
5.3 Reward Design for Online Learning
In this section, we aim to analyze the contribution of different components of the reward function as defined in 6. We fixed $\alpha$ as 1.0 and varied the weights $\beta$, and $\gamma$.
As shown in the results in Table 7, the formatting reward $r_{\text{format}}$ is crucial for both models, as its removal leads to significant performance degradation. Conversely, the influence of the strength reward $r_{\text{strength}}$ differs between the models. For Qwen-2.5-7B-Inst, removing $r_{\text{strength}}$ results in a performance improvement, while it offers a slight enhancement for the more capable 32B model. We hypothesize that 32B model’s stronger inherent abilities allow it to effectively learn the correlation between preference strength and judgment. In contrast, for the weaker 7B model, this additional learning objective may introduce confusion.
Figure 4: Comparison of different methods on data groups with different quality. Group 1 is with lower quality and Group 4 is with higher quality.
5.4 Thinking Makes the Judge More Robust
In real applications, it is common that there exists noise in the training set, or there is a distributional difference between the training and test sets. In such cases, Think-J demonstrates superior robustness compared with classifiers as a result of its judgment thinking ability.
To verify this, we adopted the approach from (Wang et al. 2024b) and divided HH-RLHF into four groups with different data quality [^11] Data quality is indicated by the difference in scores assigned to response pairs by an external reward model. For more details please refer to the work of (Wang et al. 2024b).. We then trained judges on the data based on classifier-judge or Think-J. As shown in Figure 4, for the two groups with higher data quality, classifier-based judge achieve comparable or even better performance. However, for the groups with lower quality, the accuracy of classifier-based methods drops significantly, even falling below random guessing. In contrast, Think-J maintains relative stability, verifying its robustness to varied data quality [^12] We also present thinking trace error analysis in Appendix A.6..
6 Conclusion
In this paper, we propose Think-J to enhance generative LLM-Judges with judgment thinking optimization. Experiment results verify the effectiveness of Think-J compared with both classifier-based and generative LLM judges. With the increasing popularity of RL-based test-time scaling methods, it is crucial to develop a reliable and stable feedback system that aligns well with real-world human preferences. In future, we will continue to explore generative judges for more accurate preference modeling.
Acknowledgements
This work was supported in part by the Jiangsu Science and Technology Major Project (BG2024031) and Nanjing University AI & AI for Science Funding (2024300540).
Appendix A Additional Experiments
A.1 Margin-based Reward for GRPO
As we have explained in Section 3.2, we design the reward function as the combination of three parts: $r_{\text{accuracy}}$, $r_{\text{format}}$, $r_{\text{strength}}$. Referring to the work of Chen et al., for the purpose of more fine-grained judgment prediction, we also tested the following reward function design:
[…]
where the judgment prediction should also be formatted as:
<think> {thinking trace} </think>
Therefore, the quality scores for Response (a) and Response (b) are [[30]] and [[50]], respectively.
Method Llama-3-8B-Instruct Qwen-2.5-7B-Instruct
RewardBench RewardBench
Chat Hard Safety Reason Average Chat Hard Safety Reason Average
online (GRPO w/ margin) 96.5 66.8 88.7 73.2 81.3 96.1 72.3 83.4 76.1 82.0
online (GRPO w/o margin) 93.9 74.3 90.3 77.8 84.1 96.1 78.6 85.9 80.4 85.3
Table 9: Experiment results of the impact of $r_{margin}$ on Helpsteer2-Pref.
This design enables fine-grained absolute response scoring, which is more useful for LLM preference optimization pipelines. However, as shown in Table 9, the introduction of margin-based reward results in performance degradation. By further inspection, we revealed a consistent issue of reward hacking, where the assigned scores were always driven to extreme values (e.g., 0 or 100) to maximize the reward. Despite experimenting with various formulations of $r_{\text{margin}}$, we were unable to effectively mitigate this problem. We believe that mitigating this reward hacking for $r_{\text{margin}}$ necessitates preference data annotated with absolute scores, which we leave for future investigation.
A.2 Efficiency of Trace Clipping
As we explained in Section 3.1 and Appendix B.1, we perform trace clipping for the thinking traces generated by Deepseek-R1 to reduce inference overhead and improve traning efficiency. To verify this, we present the results of thinking trace initialization with and without trace clipping in Table 10, which clearly shows that when initialization is performed without trace clipping, the optimization effect deteriorates, particularly for online RL. This finding supports our earlier hypothesis that for general LLM judgment, longer thinking do not necessarily correlate with better accuracy.
Init RL RewardBench
Chat Safety Reason Overall
Llama-3-8B-Inst 64.8 76.5 63.4 68.8
w/o clipping offline 82.6 88.1 70.2 81.4
online 79.6 85.1 67.5 78.7
w/ clipping offline 82.8 87.4 77.2 83.2
online 82.9 90.3 77.8 84.1
Table 10: Comparison of thinking trace initialization with or without trace clipping.
Moreover, beyond these performance gains, trace clipping also significantly reduces inference time and computational requirements. To quantify this, we evaluated two online-optimized Qwen-2.5-7B-Instruct models from Table 6 on a single A100-80G GPU. This evaluation involved bidirectional inference across 5,970 samples from RewardBench, and we measured both total inference time and average thinking trace length (tokenized by the Qwen2.5-7B-Instruct tokenizer). As evident from Table 11, trace clipping drastically reduced inference time, primarily due to the considerable decrease in thinking trace length.
In conclusion, by fostering a more concise thinking process for judgment thinking initialization, trace-clipping not only accelerates inference but also enhances optimization performance.
Init RL Inference Time(s) Avg Trace Length
w/o clip online 2947.01 907.06
w/ clip online 785.42 231.26
Table 11: Comparison of inference efficiency of trace clipping on Qwen-2.5-7B-Instruct.
A.3 Critic-model for Preference Pair Construction
While some of previous construct preference pairs for DPO by sampling multiple responses, In Section 3.2, we primarily employ a critic model for constructing the preference pairs. This is because, after judgment thinking initialization, the model already possesses strong judging capabilities. This means simply sampling multiple times often doesn’t guarantee the creation of distinct negative examples for much of the preference data.
To verify this, we compared both the efficiency and the performance of our critic-based method and the sampling-based approach on Llama-3-8B-Instruct. Specifically, for each instruction, we sampled 16 judgments, randomly selecting one correct judgment as the positive sample and one incorrect judgment as the negative sample. If a pair of both correct and incorrect judgments couldn’t be found, the sample was dropped.
Method Original Sampling-based Critic-based
Pair Num 9125 5111 8942
Table 12: Comparison of preference sample number created by different methods on Llama-3-8B-Instruct.
Method RewardBench
Chat Hard Safety Reasoning Overall
Baseline 90.4 44.7 76.5 63.4 68.6
Sampling-based DPO 92.9 75.0 86.9 70.1 81.2
Critic-guided DPO 95.0 73.3 87.4 77.2 83.2
Table 13: Comparison of offline learning from sampling-based or critic-guided preference pairs on Llama-3-8B-Instruct.
Data RewardBench
Chat Hard Safety Reasoning Overall
Hybirated 95.0 73.3 87.4 77.2 83.2
Solely-Critic 94.9 72.6 88.1 77.6 83.3
Table 14: Comparison of training of on data constructed solely from the critic model or in a hybrid way on Llama-3-8B-Instruct.
As shown in Table 12, for the Helpsteer2-Pref dataset, even with n=16 samples, the sampling-based method failed to generate negative samples for nearly half of the data. This significantly reduces data utilization efficiency. Therefore, our critic-based method substantially outperforms the sampling-based method, as demonstrated in Table 13.
Moreover, as we opt to using the critic model to only generate the contrast samples for the preference pairs, we also verified the effectiveness of training solely on data generated by the critic. As shown in Table 14, there was no significant difference in performance. Therefore, we decided to use the critic to construct either positive or negative samples for a given response. This approach effectively reduces inference time by half compared to constructing both.
A.4 Statistical Indicators of Judgment Thinking Optimization
In this section, we present the detailed statistical indicators during Judgment Thinking Optimization. The statistical indicators of offline learning, GRPO training, PPO (Schulman et al. 2017) training and Reinforce++ (Hu et al. 2025) training are presented in Figure 5, 6, 8, 7, respectively.
As the results indicate, GRPO demonstrates consistent and substantial increases in predicted rewards, a strong signal of learning to generate desirable outputs, and the critic’s low and stable value function loss suggests reliable outcome evaluation. While DPO effectively aligns with preferences, and Reinforce++ and PPO show progress in reward acquisition and stability, GRPO’s robust reward improvement combined with a well-functioning critic positions it as the most robust algorithm based on these indicators. Notably, GRPO’s response length exhibits a consistent gradual increase, similar to prior work of (Shao et al. 2024), while other methods start to fluctuate after a certain point. This suggests GRPO effectively learns to generate more detailed reasoning, enhancing judgment accuracy.
Figure 5: The variation of statistical metrics during offline learning trained on Helpsteer2-Pref.
Figure 6: The variation of statistical metrics during GRPO training on Helpsteer2-Pref.
Figure 7: The variation of statistical metrics during Reinforce++ training on Helpsteer2-Pref.
Figure 8: The variation of statistical metrics during PPO training on Helpsteer2-Pref.
A.5 Human Evaluation of Thinking Trace Interpretability
To quantitatively verify the improvement of thinking trace interpretability introduced by our method, we conducted a human evaluation. We randomly sampled 100 cases from RewardBench and enlisted three graduate students to manually assess the thinking trace produced by three judge models: the original model, the model fine-tuned with Supervised Fine-Tuning (SFT) on positive data only, and our model optimized with online Reinforcement Learning (RL), Think-J. The students were instructed to compare the two judgments given by different models for the same sample and determine which one provided a better thinking. We then recorded the majority vote from the three students.
Setting Win Rate
Think-J v.s. Original Qwen-2.5-7B-Inst 100%
Think-J v.s. SFT on Positive Samples 67%
Table 15: Comparison of interpretability of thinking traces generated by different judges.
As the results show in Table 15, Think-J unquestionably outperforms the judgments generated by the original model. This is expected, as the original model was not optimized for thinking-enhanced judgment, and its interpretations were too simplistic. Furthermore, Think-J also surpasses the judgments from the model SFT on positive samples only, underscoring the necessity of negative samples when optimizing judgment thinking traces.
A.6 Limited Failure Analysis of Think-J Robustness
While we generally assume a correct thinking trace would lead to a correct judgment (and vice versa), we acknowledge cases where an incorrect thinking trace might still produce the right outcome. To perform a more in-depth evaluation of the model’s underlying thinking process, we leveraged an LLM-as-a-Judge approach based on the data and model in Section 5.4.
Model Error Rate
Chat Hard Safety Reasoning
Think-J-Qwen-2.5 trained on Group1 32.5 55.0 60.0 63.1
Think-J-Qwen-2.5 trained on Group4 23.1 33.3 46.7 35.0
Table 16: Error rate of thinking traces generated by different Think-J models trained on different qualities groups of data.
Specifically, we randomly selected 50 samples from RewardBench and employed GPT-4o to meticulously check for errors in the judgment process (i.e., the thinking trace) for each model. We removed the final binary judgment (i.e., ”Response A/B is better”) and focused solely on verifying errors within the clipped critique itself, even when the final judgment result was correct. This methodology allowed us to better assess the robustness of Think-J’s internal logic.
Unexpectedly, as shown in Table 16, our analysis revealed a significant finding: Despite the high accuracy of the model trained on noised Group 1 data, its thinking trace quality appeared to degrade substantially, with over 50% of the traces containing identifiable errors. We hypothesize that for a binary classification task like judgment, an imperfect thinking trace can still lead to a correct judgment in many cases, provided the critical point with decisive influence is identified.
Appendix B Implementation Details
B.1 Thinking Trace Clipping
As we have explained in Section 3.1, despite the superior reasoning capability of Deepseek-R1, the thinking trace generated are excessively long and could possibly cause difficulty for further optimization and increase computational overhead. Specifically, for our case, we hypothesize that general LLM judgment does not necessarily require a detailed, step-by-step reasoning process. On the contrary, identifying one key factor is sufficient to make correct judgment in most cases. Overly long reasoning traces can cause the model to consider factors that do not affect the preference outcomes, thereby degrading the effectiveness.
Therefore, we introduce Trace Clipping to enhance the effectiveness of thinking initialization, as shown in the Example of 9. Specifically, we observed that the CoTs generated by R1 are composed of two parts: a lengthy reasoning process, and an explanation which summarizes the reasoning process. Therefore, we remove the first part and use the second part as the output trace. Models trained with the clipped trace will also generate the reasoning trace in a way that directly addresses the key points. This not only facilitates subsequent optimization but also enhances the judgment readability.
The experiment analysis of the effectiveness of thinking trace clipping is presented in Appendix A.2.
Figure 9: The illustration of Trace Clipping. The first part of thinking trace is removed, while the second part is used as the output thinking trace.
B.2 Judgment Thinking Initialization
As we have explained in Section 3.1, we apply deduplication to ensure the diversity of LIMJ707. For this purpose, we leverage the algorithm defined in Algorithm 1 to create instruction data pool, where the embeddings of instructions are derived based on the model of instructor-large [^13] https://huggingface.co/hkunlp/instructor-large, and repetitive instructions are removed based on the similarity of instruction semantics.
We leverage the prompt templates in Figure 11 and 13 for constructing thinking trace based on proprietary thinking models. The temperature is set as 0.0 to enable more accurate judgment.
Algorithm 1 Diversity-verified Sampling
0: Instruction Dataset $D$ with $m$ samples, number of samples to select $n$.
0: Selected Dataset $D^{\prime}$ with $n$ samples.
1: Derive the embeddings for each sample in $D$.
2: Random sample one data point $x$ from $D$.
3: Delete $x$ from $D$, add $x$ to $D^{\prime}$.
4:for i=1,2,,n1i=1,2,...,n-1 do
5:  Calculate the cosine similarity score between $x$ and each sample from $D$.
6:  Select the least similar sample $x^{\prime}$ from $D^{\prime}$.
7:  Let $x=$ $x^{\prime}$.
8:  Delete $x$ from $D$, add $x$ to $D^{\prime}$.
9:end for
Figure 10: The sampling algorithm for LIMJ707 used for diversity verification.
B.3 Judgment Thinking Optimization
For judgment thinking optimization, we leverage the prompts in Figure 11 and 12 for offline learning, and the prompts in Figure 13 and 14 for online learning. The offline learning is implemented based on Llama-Factory (Zheng et al. 2024), and online learning is implemented based on verl (Sheng et al. 2024). We present the hyper-parameter settings in Table 17 and 18.
B.4 Criteria for Determining Preference Strength
As we have explained in Section 3.2, for Rule-based Online Learning, we propose an $r_{strength}$ for assessing preference strength, which is defined as the degree to which the judge favors one response over another. In this section, we would like to detail how we get the strength annotation for HelpSteer2-Pref and HH-RLHF datasets.
For the HelpSteer2 dataset, we directly adopt the original data annotations, which were derived from professional human annotators and provide three distinct strength levels.
For the HH-RLHF dataset, we derive strength annotations following the method of (Wang et al. 2024b), which involves:
  1. 1.
    Training multiple Bradley-Terry classifiers on preference data.
  2. 2.
    Averaging the scores assigned to each data point by these classifiers.
  3. 3.
    Categorizing the strength into three tiers (1, 2, 3) based on the score differences between preference pairs.
We opted for a three-level categorization of strength—representing ”slightly better,” ”better,” and ”much better”—due to its clear interpretability. We found that further granularity did not improve results and, in some cases, led to the model ”hacking” to extreme values, as detailed in Appendix B.1.
hyper-parameter SFT offline BT classifier
learning_rate 5.00E-06 1.00E-06 2.00E-06
per_device_train_batch_size 64 64 64
cutoff_len 4096 4096 4096
lr_scheduler cosine cosine cosine
train_epoch 3 2 2
kl_beta 0.1
deepspeed_stage zero3 zero2
precision bf16 bf16 bf16
flash_attn fa2 fa2
Table 17: Hyper-parameter settings for SFT, offline learning and BT classifier.
hyper-parameter GRPO PPO Reinforce++
data.train_batch_size 1024 1024 1024
data.max_prompt_length 1024 1024 1024
data.max_response_length 3072 3072 3072
actor_rollout_ref.actor.optim.lr 1.00E-06 1.00E-06 3.00E-06
actor_rollout_ref.actor.ppo_mini_batch_size 256 256 1024
actor_rollout_ref.actor.ppo_micro_batch_size_per_gpu 16 4 8
actor_rollout_ref.actor.use_kl_loss true true
actor_rollout_ref.actor.kl_loss_coef 0.001 0.001
actor_rollout_ref.actor.kl_loss_type low_var_kl mse
actor_rollout_ref.rollout.log_prob_micro_batch_size_per_gpu 16 4 8
actor_rollout_ref.rollout.n 8 4
actor_rollout_ref.ref.log_prob_micro_batch_size_per_gpu 16 4 8
critic.optim.lr 1.00E-05
critic.ppo_micro_batch_size_per_gpu 4
algorithm.kl_ctrl.kl_coef 0.001 0.001
trainer.n_gpus_per_node 8 8 8
trainer.nnodes 2 2 2
trainer.total_epochs 10 5 5
Table 18: Hyper-parameter settings for online learning algorithms.
Figure 11: The prompt template used for judgment prediction without $r_{strength}$.
Figure 12: The prompt template used for critique prediction without $r_{strength}$.
Figure 13: The prompt template used for judgment prediction with $r_{strength}$.
Figure 14: The prompt template used for critique prediction with $r_{strength}$.